OpenStudy (superhelp101):

The reaction shown below has a negative enthalpy change and a positive entropy change. 2O3 (g) yields 3O2 (g) Which of the following best describes this reaction? The reaction will only be spontaneous at high temperatures. The reaction will only be spontaneous at low temperatures. The reaction will not be spontaneous at any temperature. The reaction will be spontaneous at any temperature.

OpenStudy (superhelp101):

@thomaster could you help me determine this?

OpenStudy (superhelp101):

@aaronq

OpenStudy (superhelp101):

@miraclemilktea2 could you help me?

OpenStudy (superhelp101):

@Abhisar

OpenStudy (superhelp101):

@iPwnBunnies

OpenStudy (aaronq):

Do you know how spontaneity is determined with these thermodynamic values?

OpenStudy (superhelp101):

no i dont

OpenStudy (aaronq):

the Gibbs free energy change (\(\Delta G\)) is indicative of the energy potentials in a dynamic equilibrium. The expression relating the enthalpy and entropy is: \(\Delta G=\Delta H-T\Delta S\) \(\Delta G<0\) reaction is spontaneous \(\Delta G>0\) reaction is non-spontaneous What you want is to know when the reaction will be spontaneous.

OpenStudy (superhelp101):

oh alright. what should i plug into the equation in order to determine it?

OpenStudy (aaronq):

yep

OpenStudy (superhelp101):

so what do i plug in?

OpenStudy (aaronq):

positive and negative signs

OpenStudy (superhelp101):

oh ok

OpenStudy (aaronq):

The question is asking at what temperatures will \(\Delta G\) be negative.

OpenStudy (superhelp101):

so would the reaction be spontaneous

OpenStudy (aaronq):

at what temperatures?

OpenStudy (superhelp101):

how do i know that?

OpenStudy (aaronq):

\(ΔG=ΔH−TΔS\) \(ΔG=-X−T(+Y)=-X-TY\) will \(\Delta G\) ever be positive? Note that T can only take on positive values (absolute scale)

OpenStudy (superhelp101):

no it can't be positive

OpenStudy (aaronq):

exactly. So \(ΔG<0\) at all temperatures.

OpenStudy (aaronq):

which means that it's always spontaneous

OpenStudy (superhelp101):

that makes sense that you very much! :)

OpenStudy (superhelp101):

thank*

OpenStudy (aaronq):

i'm glad it does! no problem !